Assuming the formation of an ideal solution, determine the boiling point of mixture containing 1560 g benzene (molar mass78)and 1125 g chlorobenzene using graph against external pressure of 1000torr
1
Expert's answer
2020-04-18T06:55:44-0400

1560 g benzene

n = m/M = 1560 / 78.11 = 19.97 mol

1125 g chlorobenzene

n = m/M = 1125 / 112.56 = 9.99 mol

Total moles:

19.97 mol + 9.99 mol = 29.96 mol

X (C6H6) = 19.97 / 29.96 = 0.6666 or 2/3

X (C6H5Cl) = 9.99 / 29.96 = 0.3334 or 1/3

The graph:



At 1000 torr and composition 2/3 benzene 1/3 Cl-benzene t = 1000C